in a shipment of fluorescent light bulbs, 1/50 of them are defective. What is the ratio of defective to non-defective light bulbs
Answer by TimothyLamb(4379)

Bad/Good = (1/50)/(49/50)
Bad/Good = (50/50)/49
---
answer:
Bad/Good = 1/49
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php